The slope of AB is 2/5. A(3, k) is a point on the line AB.  Find k if B(-7, 0).
Answer by josgarithmetic(39617)

Form the expression for slope using the given coordinates for A and B, and you already know this slope must be 2/5. Find k.

(k-0)/(3-(-7))=2/5
k/10=2/5
k=20/5
k=4